### v3

# AI as Coordination-Compressing Capital:

Abstract
Task-based models of AI and labor hold organizational structure fixed. We introduce agent capital: AI that reduces coordination costs, expanding spans of control and enabling endogenous task creation. Five propositions characterize how coordination compression affects output, hierarchy, manager demand, wage dispersion, and the task frontier. The model generates a regime fork: the same technology produces broad-based gains or superstar concentration depending on who benefits from coordination compression. Simulations with heterogeneous workers confirm sharp regime divergence. Within the simulated firm, inequality falls in all regimes as coordination compression expands employment, but the manager–worker wage gap widens universally. The distributional impact hinges on who controls organizational elasticity.
JEL Codes: D23, J24, L23, O33
Keywords: Artificial intelligence, coordination costs, organizational design, span of control, task-based models, inequality
1 Introduction
The dominant framework for analyzing AI’s labor market effects is task-based: technology changes which tasks are feasible, who performs them, and at what cost. Acemoglu and Restrepo (2019) formalize this as a displacement-reinstatement dynamic, where automation removes workers from existing tasks while new task creation reinstates labor demand. Agrawal et al. (2026) refine this approach, showing that AI can enhance worker productivity without automating tasks—by expanding the technologically feasible task set ($\bar{\tau}$) or broadening individual workers’ feasible task sets ($\Omega(i)$). Their key insight is that productivity (captured by geometric-mean uniformity of the task distribution) and inequality (captured by Shannon entropy, which measures distributional uniformity and thus moves inversely with inequality) are different functionals of the same equilibrium distribution and can move independently or even in opposite directions.
These models share a structural limitation: they hold organizational architecture fixed. Workers are reallocated across tasks, but the hierarchy through which work is coordinated—spans of control, management layers, reporting structures—is treated as exogenous. This matters because a growing body of evidence suggests that AI’s organizational effects may be as consequential as its task-level effects. Ewens and Giroud (2025) document that U.S. public firms flattened their hierarchies following AI adoption, reducing management layers across a sample of more than 3,100 firms. Babina et al. (2025) find a parallel pattern: firms investing in AI shifted toward flatter structures with proportionally fewer mid- and senior-level employees. Bloom et al. (2014) show that information technology and communication technology have distinct and sometimes opposing effects on firm organization—information technology (IT) decentralizes while communication technology (CT) centralizes—suggesting that the organizational channel deserves explicit modeling. Bresnahan et al. (2002) find that IT investment raises productivity only when accompanied by complementary organizational restructuring, implying that the reorganization channel may dominate the direct task-level channel. A related limitation concerns measurement: standard growth accounting absorbs AI’s contribution into TFP, obscuring its role as a distinct input rather than an augmentation of existing capital or labor (Farach et al., 2025).
Recent papers embed AI in Garicano’s (2000) knowledge hierarchy but hold organizational costs fixed. Ide and Talamás (2025) show that autonomous AI benefits the most knowledgeable while non-autonomous AI benefits the least knowledgeable, but their communication cost parameter is exogenous—AI reshuffles occupational roles within an unchanged organizational architecture. Xu et al. (2025) show that generative AI can paradoxically narrow spans of control through induced deskilling, but again hold communication costs fixed while varying the knowledge partition between workers and managers. Cheng et al. (2025) model AI replacement in teams with sequential information flow, showing that positional structure—not just skill level—determines displacement risk. Our approach is complementary: we hold the knowledge partition fixed and vary the coordination cost itself. This is precisely the channel that Garicano and Rossi-Hansberg (2006) showed has first-order effects on inequality—reductions in communication costs widen spans, compress hierarchies, and concentrate returns at the top—but that neither Ide and Talamás (2025) nor Xu et al. (2025) parameterizes as a function of AI capability.
This paper makes three contributions. First, we introduce agent capital ($K_{A}$) as a distinct production input that reduces coordination costs within organizations. Unlike task-level AI, which changes what workers can do, agent capital changes how work is organized—compressing the coordination friction $c(K_{A})$ that determines spans of control, hierarchy depth, and the boundary between managerial and production labor. Second, we derive five formal propositions characterizing the comparative statics of coordination compression: output rises (Proposition 1), spans of control expand (Proposition 2), manager demand falls (Proposition 3), wage dispersion rises under elite complementarity (Proposition 4), and the task frontier expands endogenously (Proposition 5). Third, these propositions generate a regime fork. When agent capital is broadly accessible (low $\beta$), coordination compression produces broad-based productivity gains and, within the firm, reduces inequality through employment expansion—a “general infrastructure” equilibrium. When agent capital disproportionately amplifies high-skill managers (high $\beta$), the same technology produces superstar concentration—an “elite complementarity” equilibrium. The distributional consequences of AI depend critically on which regime obtains.
We develop this framework using a newsroom as a motivating illustration: an Editor-in-Chief, Section Editors, and Reporters organized along a task continuum from routine briefs to elite editorial strategy. This setting grounds the model’s abstractions in a concrete organizational context—coordination friction, span constraints, and the tension between hierarchy and efficiency—and motivates the regime fork through the divergent ways coordination compression can reshape a single organization.
The paper proceeds as follows. Section 2 presents the model. Section 3 states five propositions with proofs. Section 4 introduces the regime fork and reports simulation evidence across four parameter regimes. Section 5 discusses testable predictions, policy implications, and limitations. Section 6 concludes.
2 Model
2.1 Production and Coordination
Consider a hierarchical firm with production workers and managers. To fix ideas, aggregate output takes the familiar Cobb–Douglas form $Y=A\cdot L_{\text{eff}}^{\alpha}\cdot M^{1-\alpha}$, where $L_{\text{eff}}$ is effective labor, $M$ is management, and $\alpha\in(0,1)$ is the labor share. We use this aggregate expression as a motivating framework; the formal model operates at the team level.
Each manager $i$ oversees a team of $n_{i}$ workers. Team output is:
$Y_{i}=A\cdot L_{\text{eff},i}^{\alpha}\cdot m_{i}^{1-\alpha},\qquad m_{i}=1$ (1)
where $A>0$ is total factor productivity and $m_{i}=1$ reflects that each team has exactly one manager. Since $m_{i}=1$, team output simplifies to $Y_{i}=A\cdot L_{\text{eff},i}^{\alpha}$, but the two-input structure motivates the wage-sharing rule: the manager receives the $(1-\alpha)$ share of team output as a residual claimant, and workers share the $\alpha$ share (see below). Aggregate output is $Y=\sum_{i}Y_{i}$. The term $L_{\text{eff},i}$ represents effective labor: raw labor discounted by coordination friction. As teams grow, coordination overhead increases—each additional worker requires managerial attention proportional to the per-worker coordination cost $c$. This captures the empirical regularity that large teams face diminishing returns from coordination losses (Bolton and Dewatripont, 1994).
Agent capital $K_{A}$ reduces coordination costs:
$c(K_{A})=\frac{c_{0}}{1+\gamma\cdot K_{A}}$ (2)
where $c_{0}>0$ is the baseline coordination friction and $\gamma>0$ is the coordination compression parameter—the effectiveness of agent capital at reducing friction. When $K_{A}=0$, coordination cost is $c_{0}$. As $K_{A}\to\infty$, coordination cost approaches zero. We model $K_{A}$ as a distinct factor rather than augmenting $K$ or $L$ because digital labor’s defining properties—scalability, non-rivalry, and context-dependent substitutability with human labor (Farach et al., 2025)—make it irreducible to either traditional capital (which depreciates predictably and is rivalrous) or human labor (which cannot be replicated at near-zero marginal cost).
When workers are heterogeneous—with individual skills $q_{j}$ varying across the task continuum, as in Agrawal et al.’s (2026) $s(\tau)$ framework—effective labor for manager $i$’s team is:
$L_{\text{eff},i}=\frac{Q_{i}}{1+c_{i}\cdot n_{i}}$ (3)
where $Q_{i}=\sum_{j\in\text{team}_{i}}q_{j}$ is the sum of worker skills on team $i$ (team quality) and $n_{i}$ is the team’s headcount. Note that $Q_{i}$ appears only in the numerator while $n_{i}$ appears only in the coordination penalty $1+c_{i}n_{i}$ in the denominator: coordination friction depends on the number of communication links—a function of headcount—not on worker skill. Quality enters through the numerator: a team of 10 skilled workers produces more effective labor than a team of 10 unskilled workers facing the same coordination overhead. The homogeneous case (all $q_{j}=1$) reduces to the baseline $L/(1+cL)$.
The two-input structure (1) motivates a Cobb–Douglas sharing rule: the manager receives $(1-\alpha)Y_{i}$ as a residual claimant, and the worker share $\alpha Y_{i}$ is divided among team members in proportion to their skill contribution, so worker $j$ on team $i$ receives $(q_{j}/Q_{i})\cdot\alpha\cdot Y_{i}$. Since $m_{i}=1$ is a discrete constraint (not a variable input), this is a rent-sharing convention, not competitive factor pricing; we adopt it for tractability and because the $(1-\alpha,\alpha)$ split is the natural division under Cobb–Douglas structure.
We present the model in terms of human managers coordinating human workers—the setting for which the richest organizational theory exists (Garicano, 2000; Garicano and Rossi-Hansberg, 2006). But the coordination-cost structure applies more broadly. A worker orchestrating a team of AI agents faces the same friction: each additional agent requires prompting, monitoring, and integration, with overhead proportional to the number of concurrent workstreams. In this interpretation, $K_{A}$ reduces the cost of agent orchestration rather than human management, $s_{i}$ indexes the worker’s coordination capability rather than managerial skill, and “span of control” generalizes to the number of concurrent processes—human or digital—a decision-maker can effectively direct. The propositions and regime fork hold in either interpretation; we develop the human-hierarchy case because it connects to existing empirical evidence (Ewens and Giroud, 2025) and organizational theory.
Notation.
Throughout, $s_{i}$ denotes managerial skill and $q_{j}$ denotes worker skill. The index $i$ refers to managers, $j$ to workers.
2.2 Span of Control
The coordination cost function implies a natural coordination capacity. We define:
$S_{i}(K_{A})=\frac{1}{c_{i}(K_{A})}$ (4)
as the maximum team size manager $i$ can effectively coordinate, given the technology of coordination. This is a technological parameter—a capacity constraint—not the solution to a manager’s optimization problem. Two closures are possible: one could treat team size as an interior choice where the manager trades off the marginal worker’s quality contribution against coordination drag in $L_{\text{eff}}$, or one could impose the capacity constraint $n_{i}\leq S_{i}$ as binding. We adopt the latter throughout. Specifically, we assume labor is sufficiently abundant that each manager’s team reaches the capacity constraint $S_{i}$. This holds when the total worker pool $N$ exceeds total supervisory demand $\sum S_{i}$, or equivalently when outside options are low enough that workers prefer team assignment at the prevailing wage share. Under this closure, $L_{\text{eff},i}$ in (3) gives team output conditional on $n_{i}=S_{i}$; spans are set by coordination technology, not by an interior optimality condition. Note an important implication: when $n_{i}=S_{i}=1/c_{i}$, the denominator of (3) becomes $1+c_{i}/c_{i}=2$, so $L_{\text{eff},i}=Q_{i}/2$ regardless of $c_{i}$. At the binding capacity constraint, coordination compression raises output not by reducing friction for a fixed team, but by expanding the team itself: lower $c_{i}$ increases $S_{i}$, admitting more workers, which raises $Q_{i}$. Proposition 1’s “holding team allocations fixed” qualifier is essential—it establishes the output effect for a given team size, while the capacity-constraint closure channels the practical effect through team expansion. [^1] An equivalent micro-foundation: each coordination link consumes $c_{i}$ units of the manager’s unit time endowment, yielding the constraint $c_{i}\cdot n_{i}\leq 1$, i.e., $n_{i}\leq 1/c_{i}=S_{i}$. This capacity constraint, rather than diminishing returns in $L_{\text{eff}}$, is what bounds team size. In the homogeneous case, $S=(1+\gamma K_{A})/c_{0}$, which is linear in $K_{A}$ and increasing in $\gamma$. In the newsroom: at baseline ($K_{A}=0$), a section editor with $c_{0}=0.3$ can coordinate $S=1/0.3\approx 3.3$ reporters. At $K_{A}=5$ with $\gamma=1$, capacity rises to $S=(1+5)/0.3=20$. The editor-in-chief can now oversee the entire reporting staff directly—section editors become redundant.
In the homogeneous case, the number of hierarchical layers required for $N$ workers is:
$\text{Layers}(N,S)=\lceil\log(N)/\log(S)\rceil$ (5)
As $S$ rises, layers compress. This is not task substitution—it is hierarchy substitution.
2.3 Elite Complementarity
Not all managers benefit equally from agent capital. We model heterogeneous returns through an elite complementarity parameter $\beta\geq 0$. Manager $i$ with skill $s_{i}\in(0,1]$ has effective agent capital:
$K_{\text{eff},i}=K_{A}\cdot s_{i}^{\beta}$ (6)
When $\beta$ is low (e.g., 0.2), $s_{i}^{\beta}\approx 1$ for all managers—agent capital is general infrastructure, benefiting everyone roughly equally. Note that for $\beta\in(0,1)$ and $s_{i}\in(0,1)$, the concave mapping $s_{i}\mapsto s_{i}^{\beta}$ actually compresses skill differences: a manager with $s_{i}=0.1$ and $\beta=0.2$ receives $K_{\text{eff}}=K_{A}\cdot 0.63$, well above their raw skill. This is a substantive assumption: low $\beta$ implies AI actively equalizes coordination capacity across managers. When $\beta$ is high (e.g., 3.0), $s_{i}^{\beta}$ is steeply convex—only top-skill managers extract substantial benefit. This captures the empirical distinction between AI as a broadly accessible productivity tool versus AI as a skill-amplifying complement to elite decision-makers.
Substituting (6) into (2), each manager’s coordination cost becomes:
$c_{i}(K_{A})=\frac{c_{0}}{1+\gamma\cdot K_{A}\cdot s_{i}^{\beta}}$ (7)
Equation (2) is the homogeneous special case ($\beta=0$ or identical skills). All subsequent propositions use the manager-specific $c_{i}$.
The distinction echoes Garicano and Rossi-Hansberg (2006), who show that reductions in communication costs widen inequality (enabling top managers to leverage larger spans) while reductions in knowledge-acquisition costs compress it (making workers more similar). Our $\beta$ parameter indexes which force dominates.
2.4 Endogenous Task Creation
Coordination compression may expand the feasible set of organizational tasks. Activities that were previously too costly to coordinate—newsletter verticals, podcast production, interactive graphics, real-time analytics—become viable as coordination friction falls. The empirical relevance of endogenous task creation is well documented: Autor et al. (2024) show that the majority of contemporary employment is in job categories that did not exist in 1940, with new work shifting from middle-skill to high-skill occupations after 1980. We model this as:
$T(K_{A})=T_{0}\cdot(1+\delta\cdot K_{A})$ (8)
where $T_{0}$ is the baseline task frontier and $\delta\geq 0$ is the task creation elasticity. [^2] The linear specification is chosen for analytical transparency. The employment result $\partial E/\partial K_{A}\geq 0$ (Proposition 5) requires only that $T^{\prime}(K_{A})>0$—i.e., that the task frontier is weakly expanding—and holds for any concave alternative such as $T(K_{A})=T_{0}(1+\delta\log(1+K_{A}))$. Linearity is a convenience, not a load-bearing assumption; diminishing returns to task creation would attenuate the rate of employment growth without altering the qualitative result. When $\delta=0$, the task space is fixed—coordination compression only reshuffles existing work. When $\delta>0$, new tasks absorb workers who might otherwise be displaced by span expansion, potentially offsetting the labor-demand reduction from hierarchy compression.
2.5 Connection to the Task-Based Framework
The model extends the Agrawal et al. (2026) framework by adding an organizational channel. When $K_{A}=0$ (no agent capital), organizational structure is fixed and only the task-reallocation channel operates: AI expands $\bar{\tau}$ or $\Omega(i)$, redistributing workers across tasks within a given hierarchy. The specific production structures differ—Agrawal et al. model a task continuum with comparative advantage, while we model manager-team pairs with coordination friction—but both hold organizational architecture exogenous. When $K_{A}>0$, coordination compression adds a structural channel. Agrawal et al.’s equilibrium descriptors—geometric-mean uniformity ($\exp(g)$) for productivity and Shannon entropy ($h$) for inequality—remain valid summary statistics, but they become endogenous to organizational structure rather than determined solely by worker reallocation within a fixed architecture.
Table 1: Variable Definitions
Variable Definition Source
$K_{A}$ Agent capital (coordination-compressing AI capability) Model construct
$c_{i}(K_{A})$ Coordination friction: $c_{0}/(1+\gamma\cdot K_{A}\cdot s_{i}^{\beta})$ Model construct
$\gamma$ Coordination compression effectiveness Parameter
$\beta$ Elite complementarity exponent Parameter
$\delta$ Task creation elasticity Parameter
$S_{i}(K_{A})$ Coordination capacity: $1/c_{i}(K_{A})$ Derived
$T(K_{A})$ Task frontier: $T_{0}\cdot(1+\delta\cdot K_{A})$ Derived
Figure 1: Coordination cost and span of control as functions of $K_{A}$ for varying $\gamma$. The left panel shows $c(K_{A})$ falling hyperbolically; the right panel shows span expanding linearly. Higher $\gamma$ produces faster compression and wider spans.
3 Propositions
We state five propositions characterizing the comparative statics of coordination compression. Figure 1 illustrates the core mechanism: coordination cost falls hyperbolically in $K_{A}$, expanding spans linearly. Propositions 1–2 are unconditional; Propositions 3–5 are conditional on parameter values, generating the regime fork.
Proposition 1 (Output Effect).
If $\gamma>0$, then holding team allocations fixed, $\partial Y/\partial K_{A}>0$: output is strictly increasing in agent capital.
Proof.
Fix an arbitrary worker allocation $\{n_{i},\text{team}_{i}\}$. Since $\partial c_{i}/\partial K_{A}=-c_{0}\gamma s_{i}^{\beta}/(1+\gamma K_{A}s_{i}^{\beta})^{2}<0$, coordination cost falls as $K_{A}$ rises for every manager. Effective labor rises: $\partial L_{\text{eff},i}/\partial K_{A}=-Q_{i}\cdot n_{i}\cdot(\partial c_{i}/\partial K_{A})/(1+c_{i}n_{i})^{2}>0$. Since $Y_{i}=A\cdot L_{\text{eff},i}^{\alpha}$ and $\alpha>0$, each team’s output rises, so $\partial Y/\partial K_{A}=\sum\partial Y_{i}/\partial K_{A}>0$. ∎
Remark. Under any allocation mechanism that is at least weakly improving—choosing among allocations at least as good as the status quo—the result extends to endogenous reallocation. The proportional-allocation rule used in the simulation (§ 4.4) is not globally optimal and can produce local output non-monotonicities when expanding the workforce admits low-skill workers whose coordination overhead exceeds their quality contribution, as documented in Section 4.4.
This holds regardless of $\beta$ and $\delta$. Any positive coordination compression raises output by reducing organizational friction. The mechanism is distinct from task-level productivity enhancement (the $\theta(\tau)$ channel in Agrawal et al. (2026)): agent capital does not make any individual task more productive but makes the organization of tasks more efficient.
Proposition 2 (Span Expansion).
Each manager’s coordination capacity $S_{i}(K_{A})=1/c_{i}(K_{A})$ is strictly increasing in $K_{A}$ for all $\gamma>0$.
Proof.
$S_{i}=(1+\gamma K_{A}s_{i}^{\beta})/c_{0}$, so $\partial S_{i}/\partial K_{A}=\gamma s_{i}^{\beta}/c_{0}>0$. The cross-partial $\partial^{2}S_{i}/\partial K_{A}\partial s_{i}=\gamma\beta s_{i}^{\beta-1}/c_{0}$ is positive for $\beta>0$: higher-skill managers see faster capacity growth. For $\beta>1$, this cross-partial is itself increasing in $s_{i}$—the capacity gap between top and average managers accelerates, producing the “Winner Takes All” mechanism. (For $\beta<1$, the cross-partial diverges as $s_{i}\to 0$; this is a boundary artifact of the power function and is bounded for any positive lower bound on managerial skill.) ∎
Propositions 1 and 2 are the core structural predictions: agent capital raises output and flattens hierarchy unconditionally. Ewens and Giroud (2025) and Babina et al. (2025) document exactly this pattern—firms that adopted AI shifted toward flatter hierarchies—providing the strongest available empirical support.
Proposition 3 (Manager Demand).
If $c_{0}<1$ (so that $S(0)=1/c_{0}>1$), the number of managers required to supervise $N$ workers is strictly decreasing in $K_{A}$.
Proof.
In a multi-layer hierarchy with homogeneous span $S>1$ and $L$ layers, total managers are $M=\sum_{\ell=1}^{L}N/S^{\ell}$. For large $N$ (so that $L\to\infty$), this converges to $M\approx N/(S-1)$; we use this limiting approximation for tractability. Since $S=(1+\gamma K_{A})/c_{0}$ and $c_{0}<1$ ensures $S>1$ for all $K_{A}\geq 0$, we have $\partial M/\partial K_{A}=-N\gamma/[c_{0}(S-1)^{2}]<0$. The multi-layer formula extends beyond the two-layer production model of Section 2; we include it to show that the comparative static generalizes to deeper hierarchies. In the two-layer case that the production model and simulation directly implement, the result is more immediate: total supervisory capacity $D=\sum S_{i}$ is increasing in $K_{A}$ (since each $\partial S_{i}/\partial K_{A}>0$ by Proposition 2), so fewer managers are needed to cover a given workforce. Under $\beta>1$, the reduction is compositionally concentrated: top managers absorb a larger share of workers while bottom managers become redundant. (The simulation holds the manager pool fixed at $n=20$, so Proposition 3 manifests as excess managerial capacity rather than explicit displacement.) ∎
Proposition 4 (Wage Dispersion).
Under proportional worker allocation, for any $\beta>0$, the Gini coefficient of managerial wages is strictly increasing in $K_{A}$ for $K_{A}>0$. The rate of increase is itself increasing in $\beta$.
Proof.
Under proportional worker allocation—each manager receives $n_{i}=\lambda S_{i}$ workers (for common $\lambda$) drawn from a pool with equal average skill, so that team quality satisfies $Q_{i}/Q_{j}=S_{i}/S_{j}$—the wage ratio between managers $i$ and $j$ reduces to $R=[(1+\gamma K_{A}s_{i}^{\beta})/(1+\gamma K_{A}s_{j}^{\beta})]^{\alpha}$. (Under PAM, where high-skill workers are matched to high-span managers, $Q_{i}$ may grow faster than $S_{i}$ for top managers, reinforcing the dispersion result.) For $s_{i}>s_{j}$ and $\beta>0$: $dR/dK_{A}=\alpha\cdot u^{\alpha-1}\cdot\gamma(s_{i}^{\beta}-s_{j}^{\beta})/(1+\gamma K_{A}s_{j}^{\beta})^{2}>0$, where $u=(1+\gamma K_{A}s_{i}^{\beta})/(1+\gamma K_{A}s_{j}^{\beta})$, so all pairwise wage ratios are strictly increasing in $K_{A}$. At $K_{A}=0$, all wages are equal (Gini =0=0). For $K_{A}>0$, normalize wages by the bottom manager’s wage: $\hat{w}_{i}=(S_{i}/S_{1})^{\alpha}$. Since $\hat{w}_{1}=1$ is fixed while every $\hat{w}_{i}$ for $i>1$ is strictly increasing in $K_{A}$, the Lorenz share of the bottom $k$ managers is $L_{k}=\sum_{i=1}^{k}\hat{w}_{i}/\sum_{i=1}^{n}\hat{w}_{i}$. For $k<n$, we show $dL_{k}/dK_{A}<0$. By the quotient rule, $dL_{k}/dK_{A}<0$ iff $(\sum_{i=1}^{k}\hat{w}_{i}^{\prime})(\sum_{i=1}^{n}\hat{w}_{i})<(\sum_{i=1}^{n}\hat{w}_{i}^{\prime})(\sum_{i=1}^{k}\hat{w}_{i})$, where primes denote $K_{A}$-derivatives. The pairwise ratio result (above) implies that $\hat{w}_{i}^{\prime}/\hat{w}_{i}$ is strictly increasing in $i$: managers with larger $s_{i}^{\beta}$ have strictly higher proportional wage growth rates. Since the bottom $k$ managers have the lowest growth rates, the weighted average $\sum_{i=1}^{k}\hat{w}_{i}^{\prime}/\sum_{i=1}^{k}\hat{w}_{i}$ is strictly less than $\sum_{i=1}^{n}\hat{w}_{i}^{\prime}/\sum_{i=1}^{n}\hat{w}_{i}$, which gives the required inequality. Therefore the Lorenz curve at $K_{A}^{\prime}>K_{A}>0$ lies strictly below the Lorenz curve at $K_{A}$, and the Gini rises monotonically. For $\beta>1$, the mapping $s\mapsto s^{\beta}$ is convex, amplifying skill differences superlinearly (“Winner Takes All”); for $\beta\in(0,1)$, the mapping is concave and the Gini increase is modest (“Gentle Compression”). ∎
Remark. The wage dispersion result depends on the Cobb-Douglas rent-sharing convention ($w_{M}=(1-\alpha)Y_{i}$, worker shares proportional to quality). Under alternative wage determination (e.g., Nash bargaining over team surplus), the magnitude and potentially the direction of this result may differ. The qualitative finding that differential span expansion amplifies managerial wage heterogeneity is more general than the specific Gini characterization.
This is the formal mechanism behind the regime fork. The same technology ($K_{A}$) produces rapid or modest inequality growth among managers depending on $\beta$—superstar concentration when $\beta>1$, broad compression when $\beta<1$. A subtlety emerges at the economy-wide level: when the simulation includes heterogeneous workers (§ 4.4), the full-economy Gini—encompassing managers, workers, and the unemployed—falls in all regimes as $K_{A}$ rises, because coordination compression expands employment (Proposition 5), moving workers from zero income to positive income. The regime fork at the economy level concerns the rate of inequality reduction: economy-wide Gini falls faster under low $\beta$ than under high $\beta$. The manager–worker wage gap, however, rises in all regimes—coordination compression universally concentrates returns in the coordinating layer.
Proposition 5 (Task Frontier Expansion).
If $\delta>0$, the task frontier $T(K_{A})$ is strictly increasing in $K_{A}$, and total employment $E(K_{A})=\min\bigl(\sum_{i}S_{i}(K_{A}),\;T(K_{A}),\;N\bigr)$ is weakly increasing in $K_{A}$.
Proof.
Part 1: $dT/dK_{A}=T_{0}\delta>0$, so the frontier is strictly increasing. Part 2: Employment $E=\min(D,T,N)$ where $D=\sum S_{i}$ is total supervisory capacity. At low $K_{A}$, the economy is demand-constrained ($E=D$, increasing by Proposition 2). As $K_{A}$ rises and $\delta>0$, the frontier ceiling also rises, so $E$ increases whether the binding constraint is capacity or the frontier. Since $D$, $T$, and $N$ are continuous in $K_{A}$, $E=\min(D,T,N)$ is continuous, ensuring no downward jumps at transitions between constraint regimes. The unemployment rate u=1E/Nu=1-E/N is weakly decreasing. The parameter $\delta$ determines whether coordination compression creates fixed-pie reallocation ($\delta=0$: employment hits the ceiling $T_{0}$) or expanding-pie absorption ($\delta>0$: new tasks absorb displaced labor). We note that this employment result is, by construction, a consequence of the monotonicity assumptions on $D$ and $T$; it characterizes the model’s mechanism rather than delivering a surprising prediction. The economy-wide inequality reduction reported in Section 4.4 follows from this extensive-margin expansion (moving workers from zero to positive income) and should be interpreted as such. ∎
4 Regime Fork and Simulation Evidence
4.1 The $2\times 2$ Taxonomy
The interaction of elite complementarity ($\beta$) and endogenous task creation ($\delta$) generates four qualitatively distinct regimes:
Table 2: Four Regimes of Coordination Compression
Low $\delta$ (fixed task space) High $\delta$ (task creation)
Low $\beta$ (general infrastructure) Gentle Compression Rising Tide
High $\beta$ (elite complementarity) Winner Takes All Creative Destruction
Gentle Compression (low $\beta$, low $\delta$)
Agent capital benefits all managers roughly equally. No new tasks are created. Hierarchy compresses gently; inequality stays flat or falls slightly; output rises moderately.
Rising Tide (low $\beta$, high $\delta$)
Broad access plus new task creation. Displaced coordination workers are absorbed into new roles. Output rises substantially; inequality remains contained. This is the optimistic “democratic expansion” scenario.
Winner Takes All (high $\beta$, low $\delta$)
Agent capital amplifies top managers selectively. No new tasks absorb displaced workers. Superstar managers scale across larger spans while others stagnate. Output rises—but inequality rises sharply. This extends Rosen’s (1981, 1982) superstar mechanism with a coordination channel.
Creative Destruction (high $\beta$, high $\delta$)
New tasks appear but top talent dominates them. The simultaneous expansion and concentration produces the most volatile dynamics: high output growth with high inequality.
4.2 Simulation Setup
We simulate a 20-manager newsroom with heterogeneous workers to illustrate the propositions for one calibration. The specific magnitudes reported below are not predictions—they depend on parameter choices. The propositions establish directions for general parameter values; the simulation makes those directions tangible. Table 3 lists the simulation parameters.
Table 3: Simulation Parameters
Parameter Value Interpretation
Seed 2026 Reproducibility
AA 1.0 TFP (normalized)
$\alpha$ 0.65 Labor share in production
$c_{0}$ 0.3 Baseline coordination friction
$\gamma$ 1.0 Coordination compression rate
$T_{0}$ 200 Baseline task frontier
$n_{\text{managers}}$ 20 Manager pool
Manager skills $[0.05,1.0]$ evenly spaced Managerial skill heterogeneity
$N_{\text{workers}}$ 400 Worker pool
Worker skills $\text{Beta}(2,5)$ quantiles Right-skewed (mean 0.286)
Assignment PAM Best workers to best managers
Allocation Largest-remainder Discrete worker assignment
$K_{A}$ range $[0,10]$ by 0.2 Agent capital sweep
$\beta_{\text{low}}/\beta_{\text{high}}$ 0.2 / 3.0 Infrastructure / elite complementarity
$\delta_{\text{low}}/\delta_{\text{high}}$ 0.0 / 0.3 Fixed tasks / task creation
Each manager $i$ receives effective agent capital $K_{\text{eff},i}=K_{A}\cdot s_{i}^{\beta}$, determining their individual coordination cost $c_{i}=c_{0}/(1+\gamma\cdot K_{\text{eff},i})$ and ideal span $S_{i}=1/c_{i}$. Workers are allocated to managers via largest-remainder integer assignment, then sorted by positive assortative matching (PAM): the highest-skill workers go to the managers with the highest effective agent capital, following Garicano and Rossi-Hansberg (2006). At each $K_{A}$, the number of employable workers is $\min(N_{\text{workers}},T(K_{A}))$; the remainder are unemployed at wage zero.
4.3 Results
We report simulation results as structural findings about the parameter space, characterizing what the model generates across regimes rather than predicting any particular empirical setting.
The regime fork is continuous.
Sweeping across the full $(\beta,\delta)$ plane at $K_{A}=5$ confirms that the four named regimes are zones in a continuous parameter space, not discrete categories (Figure 2). The Gini surface rises steeply with $\beta$ and is relatively insensitive to $\delta$; the output surface rises with both $\beta$ and $\delta$ but more steeply with $\delta$. The distributional consequences of coordination compression depend on a continuous parameter ($\beta$), not a binary switch.
Figure 2: Regime heatmap at $K_{A}=5$ showing manager Gini (left) and output (right) across the $(\beta,\delta)$ plane. The inequality surface rises primarily along the $\beta$ axis; the output surface rises along both axes.
Economy-wide inequality reveals what the managerial layer hides.
The full-economy Gini—encompassing managers, employed workers, and the unemployed—is dramatically higher than the manager-only Gini, starting near 0.83 across all regimes at $K_{A}=0$ and declining to as low as 0.43 under Rising Tide at $K_{A}=10$, compared with manager-only Gini values of 0.07 to 0.40. Most inequality in this economy is between layers (managers vs. workers), not within the managerial layer. The regime fork holds economy-wide: inequality falls in all regimes as $K_{A}$ rises and employment expands, but falls faster under low $\beta$ (Rising Tide) than under high $\beta$ (Winner Takes All).
Figure 3: Full-economy Gini (solid) vs. manager-only Gini (dashed) across the four regimes. The gap between the lines reveals the between-layer inequality that manager-only metrics miss. The regime fork persists at the economy level: low-$\beta$ regimes show faster inequality reduction as $K_{A}$ rises.
Wage distributions reshape as coordination compresses.
Figure 4 illustrates how distributional shapes diverge across regimes. At $K_{A}=0$, workers and managers occupy overlapping wage ranges—teams are small, output is modest, and the coordination premium is limited. As $K_{A}$ rises, the distributions separate. Under Rising Tide, the worker distribution concentrates at low wages as employment expands and newly absorbed lower-skill workers earn modest incomes, while the manager distribution extends rightward. Under Winner Takes All, worker wages concentrate near zero while the top manager’s team earns substantially more—the distributional shape itself encodes the regime.
Figure 4: Wage densities for employed workers (blue) and managers (red) at four levels of $K_{A}$ (0, 3, 7, 10), comparing Rising Tide (top row) and Winner Takes All (bottom row). The progressive separation of manager and worker distributions visualizes the regime fork in distributional shape.
Six metrics, one story.
A dashboard (Figure 5) tracking output, economy-wide Gini, manager Gini, the manager–worker wage gap, top-10% earnings share, and unemployment rate across all four regimes confirms that the regime fork is robust across multiple outcome measures. Output rises in all regimes (Proposition 1). Manager Gini diverges sharply by $\beta$ (Proposition 4). Unemployment falls as the task frontier expands, especially under high $\delta$ (Proposition 5).
Figure 5: Six-panel dashboard showing output index, economy Gini, manager Gini, manager–worker wage gap, top-10% earnings share, and unemployment rate across all four regimes under PAM. Lines are distinguished by color, linetype (solid vs. dashed), and point markers at integer $K_{A}$ values. Each metric tells a consistent story; taken together, they confirm that the regime fork is the paper’s central structural finding.
4.4 Economy-Wide Distributional Results
The heterogeneous-worker extension yields three findings that the manager-only simulation cannot produce.
First, between-layer inequality dominates within-layer inequality. As Figure 3 shows, analyses focused on managerial earnings systematically understate the inequality generated by coordination compression. The manager–worker divide is the primary source of inequality, and it widens as coordination compression allows managers to leverage larger teams.
Second, the manager–worker wage gap rises in all regimes. For this calibration, the mean manager earns approximately $1.8\times$ the mean employed worker at $K_{A}=0$, rising to approximately $5.4\times$ at $K_{A}=10$ under Gentle Compression. The direction is not regime-dependent: coordination compression universally concentrates returns in the coordinating layer, because managers capture a fixed share of team output ($(1-\alpha)Y_{i}$) while the worker share ($\alpha Y_{i}$) is divided among all team members. As spans expand, each manager’s output grows while workers share among more colleagues. This result is a direct consequence of the Cobb–Douglas rent-sharing convention: the manager captures $(1-\alpha)$ of a growing pie while the worker share is diluted across an expanding team. Under alternative wage determination (e.g., Nash bargaining or competitive labor markets with outside options), the gap trajectory could differ. The qualitative mechanism—that coordination compression amplifies the organizational leverage of the coordinating layer—is more general than the specific magnitude.
Third, employment expansion is the primary equalizing force. At $K_{A}=0$, only approximately 67 of 400 workers are employed (demand =20=20 managers $\times$ span $\approx 3.3$). As $K_{A}$ rises, spans expand and—especially when $\delta>0$—the task frontier widens, absorbing previously unemployed workers at positive wages. This is Proposition 5 made concrete: the economy-wide Gini falls primarily because workers move from zero income to positive income, not because employed workers’ wages converge. Under Rising Tide (low $\beta$, high $\delta$), the unemployment rate falls from over 80% to near zero; under Winner Takes All (high $\beta$, low $\delta$), the decline is slower because the fixed task frontier limits absorption.
A note on Proposition 1 and the simulation: while the proof establishes $\partial Y/\partial K_{A}>0$ holding team allocations fixed, the simulation uses proportional allocation with an expanding workforce. When the task frontier grows ($\delta>0$), newly admitted low-skill workers can temporarily add more coordination friction than quality, producing local output decreases along the $K_{A}$ path. This is a feature of the allocation rule, not a violation of the proposition: at any fixed $K_{A}$, output is higher than it would be at lower $K_{A}$ with the same workforce.
In an economy at full employment, the extensive margin disappears and the regime fork’s distributional consequences would operate entirely through the intensive margin—wage dispersion among the employed—which rises in all regimes. The finding that economy-wide Gini falls is specific to settings where coordination compression expands employment substantially.
Results are qualitatively robust to random (non-assortative) worker assignment: all five qualitative checks (output rising, manager Gini ordering, economy-wide Gini falling, gap rising, employment ordering) pass under random matching for all four regimes. PAM is the primary specification because it arises as an equilibrium in knowledge hierarchies where higher-skill managers generate larger team surplus and can bid more for higher-skill workers (Garicano and Rossi-Hansberg, 2006). In the present model, PAM is imposed rather than derived from a competitive assignment game, and the regime fork’s qualitative structure does not depend on it. PAM sharpens the magnitudes—particularly the Gini levels under high $\beta$—but the directional results are driven by coordination compression, not the assignment rule.
4.5 Robustness to Labor Share
The simulation uses $\alpha=0.65$ as the baseline labor share. To confirm that the regime fork’s qualitative ordering is not an artifact of this choice, we re-run the simulation at $\alpha\in\{0.50,0.65,0.80\}$ and check five qualitative features at $K_{A}=10$: (1) output rises in all regimes, (2) manager Gini is higher under high $\beta$ than low $\beta$, (3) economy-wide Gini falls in all regimes, (4) the manager–worker gap rises in all regimes, and (5) employment is weakly higher under high $\delta$ than low $\delta$. Table 4 reports the results.
Table 4: Regime Fork Robustness to Labor Share ($\alpha$). [^4] Employment values are identical across $\alpha$ by construction: spans and the task frontier depend on $\gamma$, $\beta$, and $\delta$ but not on $\alpha$.
$\alpha$ Metric GC RT WTA CD Ordering holds?
0.50 Mgr Gini 0.110 0.213 0.295 0.315 WTA >> GC, CD >> RT
0.50 Gap ($\times$) 10.0 20.0 10.0 12.5 All >> baseline
0.50 Employed 200 400 200 250 RT $\geq$ GC, CD $\geq$ WTA
0.65 Mgr Gini 0.143 0.267 0.378 0.401 WTA >> GC, CD >> RT
0.65 Gap ($\times$) 5.4 10.8 5.4 6.7 All >> baseline
0.65 Employed 200 400 200 250 RT $\geq$ GC, CD $\geq$ WTA
0.80 Mgr Gini 0.175 0.316 0.454 0.479 WTA >> GC, CD >> RT
0.80 Gap ($\times$) 2.5 5.0 2.5 3.1 All >> baseline
0.80 Employed 200 400 200 250 RT $\geq$ GC, CD $\geq$ WTA
All five qualitative checks pass at every $\alpha$ value (15/15). The magnitudes shift—higher $\alpha$ compresses the manager–worker gap and raises manager Gini—but the regime fork’s qualitative ordering is invariant. This is expected: the propositions are derived for general $\alpha\in(0,1)$, so the qualitative comparative statics should survive any interior value.
5 Discussion and Policy Implications
5.1 Testable Predictions
The model generates testable organizational predictions. Propositions 23 predict declining management share and expanding spans of control as coordination-compressing AI tools diffuse—patterns documented by Ewens and Giroud’s (2025) finding that firms flattened hierarchies after AI adoption. The critical empirical question is how AI coordination adoption covaries with managerial skill, which determines $\beta$ and thus the regime. These constructs are operationalizable via O*NET task taxonomies and AI exposure indices, though causal identification of coordination compression effects remains an open design challenge.
5.2 Three Policy Frames
The coordination-compression model suggests three distinct policy frames for AI’s labor market effects, each associated with a different view of the technology:
  1. 1.
    Replacement frame $\to$ defensive redistribution (UBI, transition assistance). This addresses AI as a substitute for labor. If coordination compression eliminates middle management roles faster than new tasks absorb displaced workers (low $\delta$), redistribution may be necessary—but it treats the symptom, not the mechanism.
  2. 2.
    Tool frame $\to$ human capital investment. This addresses AI as a complement to workers, following Agrawal et al.’s (2026) emphasis on expanding $\Omega(i)$ through education and multidimensional skill development. Policy focuses on training workers to exploit the expanded task frontier. This is necessary but insufficient if organizational structure—not worker capability—is the binding constraint.
  3. 3.
    Coordination frame $\to$ access and organizational design policy. This addresses AI as coordination-compressing capital. The critical policy lever is $\beta$: who has access to coordination-compressing AI? If access is broad, the general infrastructure regime obtains—hierarchy flattens, bottlenecks clear, gains are distributed. If access is concentrated among elite managers, the superstar regime obtains—the same technology widens inequality. Policy instruments include: ensuring broad access to AI coordination platforms, antitrust attention to AI tool concentration, and organizational design standards that distribute coordination gains.
5.3 $\beta$ as the Policy Lever
The regime fork elevates $\beta$ from a model parameter to a policy variable. Unlike $\gamma$ (which depends on technological capability) or $\delta$ (which depends on the nature of work), $\beta$ is partially shaped by institutional choices: platform pricing, licensing, training access, and organizational culture. This makes the distributional consequences of AI partially subject to policy influence—a more actionable framing than “AI will displace X% of workers.”
A natural question is whether $\beta$ could be negative—whether AI might help low-skill coordinators disproportionately, leveling rather than amplifying skill differences. Empirical evidence supports this for production tasks: Brynjolfsson et al. (2025) find that generative AI narrows the performance gap between novice and expert customer service agents. But coordination tasks differ from production tasks in a critical respect. Production tasks are often codable—AI can provide expert-level scripts to novices. Coordination tasks require judgment under ambiguity: deciding what to delegate, monitoring heterogeneous outputs, resolving conflicts between workstreams. AI provides the capacity to handle more coordination links, but the quality of coordination decisions remains skill-dependent. The model restricts $\beta\geq 0$ because coordination complementarity (AI amplifies coordination skill) is more plausible than coordination substitution (AI replaces coordination judgment). If $\beta$ were negative, the mapping $s_{i}\mapsto s_{i}^{\beta}$ would invert the ranking: managers with lower raw skill would receive more effective agent capital ($s_{i}^{\beta}>1$ for $s_{i}<1$ when $\beta<0$). For managers with skills bounded away from zero, this would compress the span distribution relative to $\beta=0$, producing a leveling effect. However, for $s_{i}$ near zero the power function $s_{i}^{\beta}$ diverges, so the distributional implications of $\beta<0$ depend on the lower bound of the skill distribution—a complication the model avoids by restricting $\beta\geq 0$.
Crucially, $\beta$ operates at every level of the organization, not only the managerial layer. A frontline worker who orchestrates AI agents to manage a complex workflow—coordinating multiple tools, monitoring outputs, integrating results—is exercising the same coordination function the model attributes to managers. The policy question is not only whether elite managers get better AI, but whether the coordination capability to use AI effectively is broadly distributed or concentrated. When $\beta$ is low, coordination compression is accessible to everyone who can direct an agent; when $\beta$ is high, only those with the skill, training, or institutional permission to orchestrate complex agentic workflows capture the gains.
5.4 Limitations
This paper has several limitations. First, the production function is stylized—a two-layer firm with Cobb–Douglas structure does not capture the complexity of real organizational design. In particular, the fixed factor shares ($\alpha$, $1-\alpha$) mechanically link the manager–worker gap to team size; under a CES specification with elasticity of substitution $\sigma\neq 1$, the managerial income share could shrink as spans expand (if management and effective labor are substitutes, $\sigma>1$), potentially attenuating the gap result. Full reversal would likely require additional changes beyond the production function alone (e.g., endogenous management intensity, alternative surplus-sharing, or non-binding capacity constraints). The regime fork’s qualitative ordering is robust to $\alpha$ (Table 4), but testing robustness to $\sigma$ is an important extension. Second, the simulations are illustrative: they demonstrate that the model generates the regime fork, but calibration to specific industries or firms requires richer data than is currently available. Third, the simulation models a single multi-manager firm, not a general-equilibrium economy—inter-firm competition, labor market clearing, and endogenous entry are abstracted away. Fourth, worker skills are exogenous and fixed; in practice, AI may alter the returns to skill acquisition, endogenizing the skill distribution. Fifth, the model abstracts from firm-boundary effects (Coase, 1937; Williamson, 1979)—coordination compression may shift the make-or-buy margin in ways not captured here. Sixth, coordination cost is strictly decreasing in $K_{A}$ with no floor; in practice, verification and auditing overhead from managing large numbers of concurrent processes may impose a lower bound on effective coordination friction, attenuating span expansion at high $K_{A}$ without altering the regime fork’s qualitative structure. Seventh, the elite complementarity parameter $\beta$ performs double duty: it simultaneously governs how pre-existing skill differences translate into coordination capacity and how AI access is distributed. In particular, the formulation $K_{\text{eff},i}=K_{A}\cdot s_{i}^{\beta}$ implies that at $K_{A}=0$, all managers have identical coordination costs regardless of skill, producing a Gini of zero as a baseline normalization rather than reflecting realistic pre-existing inequality. Separating baseline managerial heterogeneity in coordination capacity from the AI complementarity rate—for instance, via $c_{i}(K_{A})=c_{0}\cdot s_{i}^{-\eta}/(1+\gamma K_{A}s_{i}^{\beta})$ where $\eta>0$ captures pre-AI skill sorting—would allow the model to distinguish between pre-existing organizational inequality and AI-induced amplification. We leave this extension for future work.
6 Conclusion
This paper extends the task-based framework for AI and labor by introducing coordination compression as a distinct channel through which agentic AI reshapes organizations. Agent capital ($K_{A}$) reduces the friction of managing workers, expanding spans of control, compressing hierarchies, and—under positive task creation elasticity—expanding the frontier of feasible work. The model generates five propositions and a regime fork: the same technology produces broad-based gains or superstar concentration depending on a single parameter ($\beta$) that indexes who benefits from coordination compression. Numerical simulations with heterogeneous managers and workers confirm sharp regime divergence across a $2\times 2$ parameter space. Within the simulated firm, the regime fork extends beyond the managerial layer: coordination compression reduces the share of unassigned workers and widens the manager–worker gap across all regimes, but the pace of inequality reduction depends critically on $\beta$. Whether these within-firm patterns aggregate to economy-wide effects depends on general-equilibrium forces—inter-firm competition, labor market clearing, and endogenous entry—that the current partial-equilibrium framework does not model.
The economic impact of AI depends not solely on the technological frontier, but on the elasticity of organizational structure in response to falling coordination costs—and on who controls that elasticity.